OpenStudy (kainui):

a+b+c+d=n For integers 0 and greater, how many solutions are there for a given n?

OpenStudy (bezerkhd):

I'm new here

OpenStudy (bezerkhd):

but..

OpenStudy (bezerkhd):

I would like to understand how to do this.

ganeshie8 (ganeshie8):

With \(n\) stars and \(3\) bars, there will be \(\dbinom{n+3}{3}\) solutions

OpenStudy (kainui):

Nice, I never learned the stars and bars way, so I had to find a different way to the answer. :P

ganeshie8 (ganeshie8):

I'd love to see how you work it w/o stars and bars...

OpenStudy (kainui):

I realized that if I took all the possible numbers and multiplied geometric series together like this, we'd have coefficients out front that count the number of ways to get that sum. \[x^a x^b x^c x^d = x^{a+b+c+d}\] So to do that, I'm basically saying \(f(n)\) is my answer: \[\sum_{n=0}^\infty f(n)x^n = \left( \sum_{n=0}^\infty x^n \right)^4\] I didn't really know how to multiply that out, seemed difficult but I figured this out pretty quick: \[ \left( \frac{1}{1-x} \right)^k = \frac{1}{(k-1)!} \frac{d^{k-1}}{dx^{k-1}}\left( \frac{1}{1-x} \right)\] So that turned this into a pretty straightforward problem, since the derivatives of polynomials is simple: \[\sum_{n=0}^\infty f(n)x^n = \frac{1}{3!} \frac{d^{3}}{dx^{3}}\left( \sum_{n=0}^\infty x^n \right) = \sum_{n=0}^\infty \frac{(n+3)(n+2)(n+1)}{6} x^n\] \[f(n) = \frac{(n+3)(n+2)(n+1)}{6} = \binom{n+3}{3}\] So pretty fun, I'm gonna try to read up on stars and bars and see if it's similar in any way to this.

OpenStudy (kainui):

I'm not gonna lie, I'm actually pretty confused by the stars and bars method, I think I drank too much coffee and can't focus on the words right now haha so I'll have to learn that way some other time I think.

ganeshie8 (ganeshie8):

Your method can be used to work these type of problems with constraints too

ganeshie8 (ganeshie8):

For example, Find the integer solutions to \(a + b + c + d = n\) where \(-2\le a\lt 5\) \(3\le b\lt 7\) \(0\le c,d\)

OpenStudy (kainui):

I didn't think of that bu now that you've got me thinking, if we have fixed constants r and s, how many solutions are there to: \(ar+bs=n\) Just amounts to replacing x with \(x^r\) and \(x^s \)in the generating function

ganeshie8 (ganeshie8):

Would that equation have solutions only if \(\gcd(r,s)\mid n\) ?

OpenStudy (kainui):

I don't know, I don't think so, I am not sure so I'll just state it as clearly as I can: r,s, and n are constant integers (let's say greater than or equal to 0 for now, not sure if this applies in the negative case) then I'm asking how many ways can I choose a and b to make this true: \[ar+bs=n\]

ganeshie8 (ganeshie8):

\[(1+x^r + x^{2r}+\cdots )(1+x^s + x^{2s}+\cdots )\] The exponents add up to \(n\) only when \(\gcd(r,s)\mid n\) right ?

ganeshie8 (ganeshie8):

so the equation \(ar+bs=n\) need not have solutions for arbitrary \(r,s,n\)

OpenStudy (kainui):

Yeah, that last thing you said is true, the gcd statement is confusing me. But yes, there's nothing wrong with having 0 coefficient since there would be 0 ways of making some combinations yes, like r=3, s=5, n= 2 for instance

ganeshie8 (ganeshie8):

It is a diophantine equation, it can have infinite "nonnegative" solutions too

ganeshie8 (ganeshie8):

r = 3, x = -5, n = 2

OpenStudy (kainui):

It can't infinite solutions, at least in the way I've phrased it here since n is fixed, it's a finite number. So if you're only adding up positive numbers you'll eventually get "too big", well hre's the generating function anyways: \[\left( \sum_{a=0}^\infty x^{ar} \right) \left( \sum_{b=0}^\infty x^{bs} \right) = \sum_{n=0}^\infty f(n) x^n \]

ganeshie8 (ganeshie8):

right, solutions will be finite if \(r\) and \(s\) are nonnegative

OpenStudy (kainui):

I just haven't really thought about it before so I'm thinking about what negative exponents will do right now trying to figure stuff out.

OpenStudy (kainui):

This is weird, I am not sure what is going on with this: \[\sum_{n=0}^\infty x^{-n} = \frac{1}{1-x^{-1}} = -x \frac{1}{1-x} = -x \sum_{n=0}^\infty x^n \]

ganeshie8 (ganeshie8):

\(\dfrac{1}{x}\lt 1 \implies x \gt 1\) so only one of the series is convergent

OpenStudy (kainui):

Heh, I wonder if that matters for formal power series though. Convergence is not necessary to use formal power series, so I wonder if this can still get useful results.

ganeshie8 (ganeshie8):

aren't we messing with below kindof stuff ? \[\dfrac{1}{1-2} = \sum\limits_{n=0}^{\infty}2^n\]

ganeshie8 (ganeshie8):

\[\sum_{n=0}^\infty 2^{-n} = \frac{1}{1-2^{-1}} = -2 \frac{1}{1-2} = -2 \sum_{n=0}^\infty 2^n\]

ganeshie8 (ganeshie8):

im not feeling comfortable haha

OpenStudy (kainui):

Hahaha well that's not true I am sure, but if you have some function attached I think it might make sense. We're no really concerned with "what's x?". I never even used x when I solved the original problem I asked, I pretty much just used it as a vehicle to combine my terms in the way I want, at no point did its convergence really matter.